• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name pking

    Log entry time 23:00:00 on October 17, 2009

    Entry number 295419

    keyword=News on Left Q3 He level; please post "tools" once an hour

    Ed has been in contact with Cryo.

    The problem does nto appear to be the level sensor. The valve controller "may" be flaky. The control loop has been modified for tonight to make it more sensitive.

    We should probably be okay for tonight. To help Ed monitor the situation, please post a shot of the tools screen once an hour.